# Probing forced responses and causal mechanisms in large-scale climate dynamics with reduced-order neural models

### v3

@@ II.2 Response theory: a (very) brief overview
Given a $n$-dimensional coarse-grained state vector $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$, we perturb a degree of freedom $x_{j}$ with a small, step function perturbation $x_{j}(0)\rightarrow x_{j}(0)+\delta x_{j}(t)$, with $\delta x_{j}(t)=\Delta_{j}\delta(t)$, where $\delta(t)$ is a delta function and $\Delta_{j}$ is a small perturbation amplitude [76, 17]. We then consider a general observable $A(x_{k}(t))$ (i.e. a general functional of the system) and quantify its response, in terms of ensemble average, as:
@@
where $\langle\cdot\rangle$ refers to ensemble average and the subscript “p” refers to the perturbed system. In this paper we are going to focus on two different specific observables: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution. The two observables allow us to capture time-dependent responses in the mean and variance of the perturbed distribution respectively. The impulse response operator $\mathbf{R}(t)\in\mathbb{R}^{n,n}$ is then defined as:
@@
The utility of the inferred impulse response operator is two-fold:
  • Causal inference. Understanding causal relations across degrees of freedoms in physical systems require us to study the effect of interventions on the system [77]. The inferred response operator quantifies time-dependent cause-effect dependencies across degrees of freedoms $x_{j}$ and $x_{k}$ [78, 79, 17, 80, 18]. We refer to Appendix A for details on the relation between the response operator and causality.
  • Response to generic time-dependent forcing. The response operator $\mathbf{R}(t)$ can be used to predict the response of the system to small, but general, time-dependent external forcing $\mathbf{f}(t)\in\mathbb{R}^{n}$ applied directly to the system’s degrees of freedoms as:
@@
A data-driven model capable of accurately capturing the impulse response operator will also capture the time-dependent causal interactions across degrees of freedom. Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, by for example assessing the causal sensitivity of the system to external perturbation [18]. Moreover, since the linear response to small perturbations is given by the convolution formula in Eq. (5), recovering the impulse response operator ensures a faithful approximation of the system’s linear response to general forcings. For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to capture the response to small, general perturbations, and is necessary (but not sufficient) to capture nonlinear responses as well. Importantly, the response operator can be inferred from data alone through the Fluctuation-Dissipation Theorem [49], this is discussed in Appendix A.
@@ III Limitations of data-driven modeling by analogy: the case of partial observations
A fundamental limitation of purely data-driven modeling of multiscale climate systems is that we cannot observe all degrees of freedom. Even when governing equations are known, as in meteorology, climate fields are only available at finite temporal and spatial resolution. Dynamics at smaller and faster scales are therefore missing from the data and often remain unparameterized in emulators, unlike in full climate models. An even greater challenge arises in many areas where we lack governing equations altogether to guide emulator construction. In both cases, data-driven approaches can mimic the observed variability of the retained variables, reproducing probability distributions, autocorrelation functions, and providing reasonable forecasts. Yet testing responses to perturbations, where unobserved dynamics becomes relevant, often exposes structural deficiencies that are otherwise hidden. The goal of this section is to illustrate such limitations within a controlled mathematical setting, where they can be made explicit and unambiguous.
@@ III.1 A triad model as a simplified prototype for coarse-grained, stochastic climate modeling
We focus on a simplified yet relevant dynamical system that retains essential features of high-dimensional geophysical flows, making the conclusions of the Section meaningful for climate applications. The model was introduced by Majda and collaborators [46] and later used in [47] in the context of climate dynamics. In what follows, we present the model and refer to Appendix B for an in-depth discussion on its relevance as a building block for complex turbulent flows. The triad model considered in this study takes the following form:
@@
This idealized system mimics the functional structure of turbulent geophysical flows, featuring linear couplings $L_{11},L_{12},L_{13}$ and energy-conserving quadratic nonlinearities. The parameter $I$ controls the strength of the nonlinearity: when $I=0$, the model reduces to purely linear dynamics and Gaussian statistics. The quadratic interactions, here between $x_{1}$ and $x_{2}$, typically arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs [16]. The parameter $\epsilon$ sets the time-scale separation between the slow mode $x_{1}$ and the fast modes $(x_{2},x_{3})$, allowing exploration of multiscale dynamics in a controlled setting. The terms $W_{2}$ and $W_{3}$ represent independent Wiener processes, introducing stochastic forcing into the second and third equations. The model generates unimodal distributions with relatively small, but significant, deviations from Gaussianity, which is a relevant feature in the study of low-frequency climate variability [81] and found in both general circulation models and observational data, see e.g. [82, 83, 84].
@@ III.2.1 Functional form of the data-driven models considered
Note that this is the only Section in this work where we compare linear and nonlinear autoregressive models. The linear method is considered for two reasons. First, we show that linear inverse modeling (LIM), as originally proposed in [3], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble mean response to external perturbations. Second, and consistent with theoretical expectations [47], we demonstrate that linear models are fundamentally limited in capturing changes in the full probability distribution: for instance, they cannot represent any response in the variance to external forcings. This limitation restricts their usefulness for studying forced changes in distributions beyond the mean.
@@ III.2.2 Stationary statistics
We integrate the two data-driven models inferred from data and specified in Eq. (7) and (8) for $T=10^{7}$, yielding time series of the same length as those obtained from the original triad system. We then compare the modeled and original probability density functions (PDFs) and autocorrelation functions (ACFs), as shown in Figure 1. The neural model described in Equation (8) successfully captures the stationary distributions of $x_{1}(t),x_{2}(t)$ and $x_{3}(t)$. In particular, it outperforms the linear model by accurately reproducing the non-Gaussian features of $x_{1}$. Both models also reproduce the autocorrelation functions well, as seen in the second row of Figure 1. In summary, the neural emulator defined in Equation (8) is capable of capturing both the PDF and autocorrelations of the full state vector $\mathbf{x}(t)$. At the same time, the linear model serves as a strong baseline for evaluating more sophisticated data-driven approaches.
@@ III.2.3 Response of probability distributions to external perturbations
We now examine the impulse response of the original triad model (Equation (6)), the linear model (Equation (7)), and the nonlinear neural model (Equation (8)). Specifically, we analyze the time-dependent response of observables $A(x_{k}(t))$, with $k=1,2,3$, to an impulse perturbation imposed on $x_{1}(0)$ at time $t=0$. We choose as observables the mean and variance of the time-dependent perturbed distribution and refer to Appendix A for details on how to compute the response operator. The results, shown in Figure 2, reveal that the nonlinear neural model captures the mean response of the distribution quantitatively, while the linear model captures it only qualitatively. The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, clearly highlighting the limitations of linear dynamics [47]. In contrast, the neural model captures the variance response with relatively high fidelity. We therefore consider the neural model presented here as a skillful model both in terms of reproducing the stationary distribution as well as the response operator.
@@
we now consider a more targeted test involving a specific, time-dependent forcing. We introduce a logarithmic forcing term to the right-hand side of the original and data-driven triad model. The forcing is applied only to $x_{1}$ and takes the form $(0.01log(1+t),0,0)$ per time step. Each model is integrated for 2500 time steps (25 Model Time Units), and we compute the difference between the forced and unforced simulations in terms of ensemble mean and variance responses. Results are presented in Figure 3. The neural model (Eq. (8)) accurately reproduces the forced responses across all three components. The linear model (Eq. (7)) performs qualitatively well but exhibits significant deviations in the ensemble mean response of the second mode $x_{2}$ as shown in Figure 3(b). This is consistent with the earlier Section: the largest error in the linear model’s impulse response also occurred in $x_{2}$ (see Figure 2(b)). This supports the interpretation that accurately capturing the impulse response operator is critical for a model’s reliability in predicting responses to general, time-dependent forcings. As expected, the linear model is not able to capture the response in ensemble variance. The neural model shows skillful response, capturing the time-dependent changes in ensemble variance relatively well, with the largest error in $x_{2}$. In general, considering that the changes in variance are very small, we conclude that the neural emulator is skillful in responding to the imposed logarithmic forcing.
@@ III.2.4 Main takeaways
Given a (practically) infinitely long trajectory of a stochastic dynamical system, and assuming the ideal condition where the full state vector is observed, it is possible to construct a neural stochastic model that is stable, reproduces the stationary statistics, and responds accurately to small but general external perturbations. The analysis does not rule out that other systems may struggle to capture responses even with full observations; however, these issues are here regarded as technical rather than fundamental. Such technical issues can be addressed by considering more complex architectures or by adding constraints withing a chosen architecture. While these results are encouraging, they remain far from any real-world scenario. In practice, we typically have access only to partial observations of a complex, interconnected system.
@@ III.3 The case of unknown equations and a partially observed state vector
We now return to the triad model defined in Eq. (6) and consider the more realistic scenario of model inference under partial observations and multiple timescales. The parameters chosen for this test are: $L_{11}=-2$, $L_{12}=0.2$, $L_{13}=0.1$, $\gamma_{2}=0.6$, $\gamma_{3}=0.4$, $\sigma_{2}=1.2$, $\sigma_{3}=0.8$, $I=1$, $\epsilon=0.01$. In this setting the variable $x_{1}$ clearly is the slow mode, forcing the much faster modes $x_{2}$ and $x_{3}$. We again integrate the system in (6) with an Euler–Maruyama scheme for a period of length $T=10^{7}$ with time step $dt=0.01$ (i.e. $10^{5}$ Model Time Units).
@@
    It must correctly capture its response to perturbations as in the original triad system.
Criterion (ii) is particularly demanding because, in the triad model defined by Eq. (6), perturbations can propagate through the system, inducing cascades of feedbacks. These time-dependent feedbacks must be effectively parametrized within the scalar model, despite the absence of direct observations of some components. While clearly nontrivial, this problem is representative of real-world data-driven modeling scenarios, where models are often built from partial observations of high-dimensional systems.
Proposing a scalar data-driven model from partial observations that satisfies the two criteria above raises two central challenges:
  • Choice of variable to model. The reduced-order model depends critically on the choice of the modeled variable. Not all components can be selected to satisfy criteria (i) and (ii) discussed above. The only viable option is to model the slow mode $x_{1}$, since the cumulative effect of the much faster $x_{2}$ and $x_{3}$ can be represented stochastically. Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model under partial observations [48, 43, 85, 86]. In realistic spatiotemporal climate dynamics, this choice translates into selecting appropriate climate fields together with a narrow range of spatial and temporal scales as we will show in Section IV.
  • The need for suitable parametrizations. Having identified the proper variable $x_{1}$ to model, the next challenge is to appropriately parametrize the cumulative effects of the unresolved components $x_{2}$ or $x_{3}$ onto the slow mode $x_{1}$. This leads to the classic problem of stochastic parametrization [87, 88, 2, 1, 67, 89]. As we will show, the success of the resulting model, particularly in capturing forced responses, hinges critically on the quality of this parametrization.
We note that for the specific case of the triad model in Eq. (6), Majda and collaborators showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ [46, 47]; see discussion in Appendix B. Our aim here is not to test such analytical strategies, which require full knowledge of the governing equations and are often limited in real-world settings; see footnote  [^1] Even moderately more complex systems already expose the practical limitations of rigorous stochastic mode reduction procedures; see Strounine et al. (2008) in the context of quasi-geostrophic modeling [8].. Rather, we use the triad model as a simple conceptual tool to highlight key challenges of data-driven modeling of multiscale systems. In realistic settings, both the governing equations and the full set of relevant variables are typically unknown [91]. Even before specifying a model structure, identifying the slow variables is nontrivial and highly problem-dependent.
@@ III.3.1 Functional form of the reduced order models considered
  • Neural model with multiplicative noise. We consider a neural model as before, but now defined with multiplicative rather than additive noise:
@@
    The multiplicative (state dependent) noise $\sigma(x)$ is fitted with an additional neural network, see Appendix C.

### v4

@@ II.2 Response theory: a (very) brief overview
Given a $n$-dimensional coarse-grained state vector $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$, we perturb a degree of freedom $x_{j}$ with a small, step function perturbation $x_{j}(0)\rightarrow x_{j}(0)+\delta x_{j}(t)$, with $\delta x_{j}(t)=\Delta_{j}\delta(t)$, where $\delta(t)$ is a delta function and $\Delta_{j}$ is a small perturbation amplitude [Risken, Baldovin]. We then consider a general observable $A(x_{k}(t))$ (i.e. a general functional of the system) and quantify its response, in terms of ensemble average, as:
@@
where $\langle\cdot\rangle$ refers to ensemble average and the subscript “p” refers to the perturbed system. In this paper we are going to focus on two observables: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution. The two observables allow us to capture time-dependent responses in the mean and variance of the perturbed distribution respectively. The impulse response operator $\mathbf{R}(t)\in\mathbb{R}^{n,n}$ is then defined as:
@@
We refer to Appendix A for details on how to infer $R_{k,j}(t)$ in practice. The utility of the inferred impulse response operator is two-fold:
  • Causal inference. Understanding causal relations across degrees of freedoms in physical systems requires to study the effect of interventions on the system [IsmaelNew]. The inferred response operator quantifies time-dependent cause-effect interactions across degrees of freedoms $x_{j}$ and $x_{k}$ [Aurell, Lucarini2018, Baldovin, Valerio2024, FabCausal]. We refer to Appendix A for details on the relation between the response operator and causality.
  • Response to generic time-dependent forcing. The response operator $\mathbf{R}(t)$ can be used to predict the response of the system to small, but general, time-dependent external forcing $\mathbf{f}(t)\in\mathbb{R}^{n}$ as:
@@
A data-driven model capable of accurately capturing the impulse response operator will also capture the time-dependent causal interactions across degrees of freedom. Response theory can then be used to assess the skills of emulators by going beyond purely statistical metrics such as r-squared, assessing the causal sensitivity of the system to external perturbation [FabCausal]. Thus, this approach opens new, rigorous avenues for evaluating and studying data-driven models. Moreover, recovering the impulse response operator ensures a faithful approximation of the system’s linear response to small general forcings. For this reason, we view the impulse response operator as a fundamental building block that a data-driven model must reproduce to capture the responses to general perturbations. Importantly, the response operator can be inferred from data alone through the Fluctuation-Dissipation Theorem [Marconi], as discussed in Section V and in Appendix A.
@@ III Limitations of data-driven modeling by analogy: the case of partial observations
A fundamental limitation of purely data-driven modeling of real-world, multiscale climate systems is that many degrees of freedom remain unobserved. Even when the governing equations are known, as in meteorology, climate fields are only sampled at finite temporal and spatial resolutions, which requires data-driven closures for unresolved fast and small-scale processes. The challenge is even greater in settings where governing equations are unknown and provide no guidance for emulator design. In both cases, data-driven approaches can reproduce the variability of the retained variables, including their probability distributions, autocorrelation functions, and short-term forecasts. However, the response of probability distributions to perturbations, particularly the mean and variance in realistic climate applications, can uncover structural deficiencies stemming from unresolved dynamics that would otherwise remain hidden from stationary statistics alone. The goal of this section is to illustrate such limitations within a controlled mathematical setting, where they can be made explicit and unambiguous.
@@ III.1 A triad model as a simplified prototype for coarse-grained, stochastic climate modeling
We focus on a simplified yet relevant dynamical system that retains essential features of high-dimensional geophysical flows, making the conclusions of this section meaningful for climate applications. The model was introduced by Majda and collaborators [NormalForms] and later used in [Majda2010] in the context of climate dynamics. In what follows, we present the model and refer to Appendix B for an in-depth discussion on its relevance as a building block for complex turbulent flows. The triad model considered in this study takes the following form:
@@
This idealized system mimics the functional structure of turbulent geophysical flows, featuring linear couplings $L_{11},L_{12},L_{13}$ and energy-conserving quadratic nonlinearities. The parameter $I$ controls the strength of the nonlinearity: when $I=0$, the model reduces to purely linear dynamics and Gaussian statistics. The quadratic interactions typically arise when geophysical turbulent flows are expressed in terms of orthogonal bases such as EOFs [majdaSIAM]. The parameter $\epsilon$ sets the time-scale separation between the slow mode $x_{1}$ and the fast modes $(x_{2},x_{3})$, allowing exploration of multiscale dynamics in a controlled setting. The terms $W_{2}$ and $W_{3}$ represent independent Wiener processes. The model generates smooth, unimodal distributions with relatively small (but significant) deviations from Gaussianity, and it exhibits fast decay of correlations. Both features are relevant in the study of low-frequency climate variability [MajdaStructuralStability] and found in both general circulation models and observational data, see e.g. [FranzkeMajda, Judith, Sura].
@@ III.2.1 Functional form of the data-driven models considered
Note that this is the only Section in this work where we compare linear and nonlinear autoregressive models. The linear method is considered for two reasons. First, we show that linear inverse modeling (LIM), as originally proposed in [Penland89], serves as a strong baseline for neural autoregressive models, successfully capturing both the mean and variance of the system’s stationary distribution, the autocorrelation functions, and the ensemble mean response to external perturbations. Second, and consistent with theoretical expectations [Majda2010], we demonstrate that linear models are fundamentally limited in capturing changes in the full probability distribution: for instance, they cannot represent any response in the variance to external forcings. This limitation restricts their usefulness for studying forced changes in distributions beyond the mean.
@@ III.2.2 Stationary statistics
We integrate the two data-driven models in Eq. (7) and Eq. (8) for $T=10^{7}$, yielding time series of the same length as those obtained from the original triad system. We then compare the modeled and original probability density functions (PDFs) and autocorrelation functions (ACFs), as shown in Figure 1. The neural model described in Eq. (8) successfully captures the stationary distribution of the triad system. In particular, it outperforms the linear model by accurately reproducing the non-Gaussian features of $x_{1}$. Both models also reproduce the autocorrelation functions well, as seen in the second row of Figure 1. In summary, the neural emulator is capable of capturing both the PDF and autocorrelations of the full state vector $\mathbf{x}(t)$. At the same time, the linear model serves as a strong baseline for evaluating more sophisticated data-driven approaches.
@@ III.2.3 Response of probability distributions to external perturbations
We now examine the impulse response of the original triad model (Eq. (6)), the linear model (Eq. (7)), and the nonlinear neural model (Eq. (8)). Specifically, we analyze the time-dependent response of observables $A(x_{k}(t))$, with $k=1,2,3$, to an impulse perturbation imposed on $x_{1}(0)$ at time $t=0$. We choose as observables the mean and variance of the distribution. The results, shown in Figure 2, reveal that the nonlinear neural model captures the mean response of the distribution quantitatively, while the linear model captures it only qualitatively. The most significant differences arise in the response of the variance: the linear model yields exactly zero variance response, as expected [Majda2010]. In contrast, the neural model captures the variance response with relatively high fidelity. We therefore consider the neural model presented here as a skillful model both in terms of reproducing the stationary distribution as well as the response operator.
@@
we now consider a more targeted test and introduce a logarithmic forcing term to the right-hand side of both the original and data-driven triad model. The forcing is applied only to $x_{1}$ and takes the form $(0.01log(1+t),0,0)$ per time step. Each model is integrated for 2500 time steps (25 Model Time Units), and we compute the difference between the forced and unforced simulations in terms of ensemble mean and variance responses. Results are presented in Figure 3. The neural model (Eq. (8)) accurately reproduces the forced responses across all three components. The linear model (Eq. (7)) performs qualitatively well but exhibits significant deviations in the ensemble mean response of the second mode $x_{2}$ as shown in Figure 3(b). This is consistent with the earlier Section: the largest error in the linear model’s impulse response also occurred in $x_{2}$ (see Figure 2(b)). This supports the interpretation that accurately capturing the impulse response operator is critical for a model’s reliability in predicting responses to general, time-dependent forcings. As expected, the linear model is not able to capture the response in ensemble variance. The neural model shows skillful response, capturing the time-dependent changes in ensemble variance relatively well, with the largest error in $x_{2}$. In general, considering that the changes in variance are very small, we conclude that the neural emulator is skillful in responding to the imposed logarithmic forcing.
@@ III.2.4 Main takeaways
Given a (practically) infinitely long trajectory of a stochastic dynamical system, and assuming the ideal condition where the full state vector is observed, it is possible to construct a neural stochastic model that is stable, reproduces the stationary statistics, and responds accurately to small but general external perturbations. The analysis does not rule out that other systems may struggle to capture responses even with full observations. These issues, while important, are considered here as technical, rather than fundamental, and point to the need for more sophisticated architectures, improved algorithms, or the inclusion of additional physical constraints within existing frameworks.
While the results shown in this Section are encouraging, they remain far from any real-world scenario. In practice, we typically have access only to partial observations of a complex, interconnected system.
@@ III.3 The case of unknown equations and a partially observed state vector
We now return to the triad model defined in Eq. (6) and consider the more realistic scenario of model inference under partial observations and multiple timescales. The parameters chosen for this test are: $L_{11}=-2$, $L_{12}=0.2$, $L_{13}=0.1$, $\gamma_{2}=0.6$, $\gamma_{3}=0.4$, $\sigma_{2}=1.2$, $\sigma_{3}=0.8$, $I=1$, $\epsilon=0.01$. In this setting the variable $x_{1}$ clearly is the slow mode, forcing the much faster modes $x_{2}$ and $x_{3}$. We again integrate the system in Eq. (6) with an Euler–Maruyama scheme for a period of length $T=10^{7}$ with time step $dt=0.01$ (i.e. $10^{5}$ Model Time Units).
@@
    It must correctly capture the mean and variance response to perturbations as in the original triad system.
Both criteria are nontrivial, with the second posing the greater difficulty. Together they mirror core challenges in data-driven modeling built from partial observations. Satisfying these criteria leads to two central challenges:
  • Choice of variable to model. The reduced-order model depends critically on the choice of the modeled variable. Not all components can be selected to satisfy criteria (i) and (ii) discussed above. The only viable option is to model the slow mode $x_{1}$, since the cumulative effect of the much faster $x_{2}$ and $x_{3}$ can be represented stochastically. Identifying such a proper variable is a necessary prerequisite for constructing a meaningful model under partial observations [MajdaBook, FutureFromPast, Lacorata, BaldovinPlos]. In realistic spatiotemporal climate dynamics, this choice translates into selecting appropriate climate fields together with a narrow range of spatial and temporal scales as we will show in Section IV.
  • The need for suitable parametrizations. Having identified the proper variable $x_{1}$ to model, the next challenge is to appropriately parametrize the cumulative effects of the unresolved components $x_{2}$ or $x_{3}$ onto the slow mode $x_{1}$. This leads to the classic problem of stochastic parametrization [ArnoldHM, Dawson, LucariniChekroun, GhilLucarini, MajdaPerspective, StochBAMS]. As we will show, the success of the resulting model, particularly in capturing forced responses in the variance, hinges critically on the quality of this parametrization.
We note that for the specific case of the triad model in Eq. (6), Majda and collaborators showed that an analytical reduced-order equation for the slow variable $x_{1}$ can be rigorously derived in the limit $\epsilon\to 0$ [NormalForms, Majda2010]; see discussion in Appendix B. Our aim here is not to test such analytical strategies, which require full knowledge of the governing equations and are often limited in real-world settings; see footnote  [^1] Even moderately more complex systems already expose the practical limitations of rigorous stochastic mode reduction procedures; see Strounine et al. (2008) in the context of quasi-geostrophic modeling [STROUNINE2010145].. Rather, we use the triad model as a simple conceptual tool to highlight key challenges of data-driven modeling of multiscale systems. In realistic settings, both the governing equations and the full set of relevant variables are typically unknown [Cecconi]: even before specifying a model structure, identifying the slow variables is nontrivial and highly problem-dependent.
@@ III.3.1 Functional form of the reduced order models considered
  • Neural model with multiplicative noise. We consider a neural model with multiplicative:
@@
    We use the same drift $(Lx+n(x))$ fitted for the additive model in Eq. (9), so that differences can only arise from the stochastic parametrization. The multiplicative (state dependent) noise $\sigma(x)$ is fitted with an additional neural network, see Appendix C.